# V2X-DGW: Domain Generalization for Multi-agent Perception under Adverse Weather Conditions

### v3

@@ Abstract
Current LiDAR-based Vehicle-to-Everything (V2X) multi-agent perception systems have shown the significant success on 3D object detection. While these models perform well in the trained clean weather, they struggle in unseen adverse weather conditions with the real-world domain gap. In this paper, we propose a domain generalization approach, named V2X-DGW, for LiDAR-based 3D object detection on multi-agent perception system under adverse weather conditions. Not only in the clean weather does our research aim to ensure favorable multi-agent performance, but also in the unseen adverse weather conditions by learning only on the clean weather data. To advance research in this area, we have simulated the impact of three prevalent adverse weather conditions on two widely-used multi-agent datasets, resulting in the creation of two novel benchmark datasets: OPV2V-w and V2XSet-w. To this end, we first introduce the Adaptive Weather Augmentation (AWA) to mimic the unseen adverse weather conditions, and then propose two alignments for generalizable representation learning: Trust-region Weather-invariant Alignment (TWA) and Agent-aware Contrastive Alignment (ACA). Extensive experimental results demonstrate that our V2X-DGW achieved improvements in the unseen adverse weather conditions.
Keywords:
3D Object Detection, Vehicle-to-Everything (V2X), Multi-agent Perception, Domain Generalization, Adverse Weather
1 Introduction
The advent of multi-agent perception systems has marked a significant leap forward in surmounting the inherent limitations of single-agent perception from the challenges of perceiving range and occlusion [3, 45]. Leveraging Vehicle-to-Everything (V2X) communication technology for the exchange of LiDAR and camera data, these systems have demonstrated a substantial enhancement in perception performance over their single-agent counterparts [39, 48]. Notably, the utilization of LiDAR sensors, preferred for their insensitivity to lighting conditions thereby facilitating effective operation both day and night, has become a cornerstone in the development of state-of-the-art cooperative perception methodologies [48, 39, 22, 3, 20]. However, these methods, predominantly reliant on supervised learning, presuppose a consistent distribution of features between training and testing datasets—a hypothesis that often fails under the real-world variability of adverse weather conditions, thus presenting a barrier to the safety-ensured autonomous driving.
Recent endeavors in Unsupervised Domain Adaptation (UDA) methods have sought to address this shortfall by feeding the unlabeled target-domain data in the model training, aiming to ameliorate performance degradation within specific target domains [15, 38, 25]. Despite their efficacy in these target domains, UDA methods might show limited generalizability to unseen data corrupted by adverse weather conditions, leading to a pronounced performance decline. This predicament stems from notable unknown domain gaps between the training data and unseen testing data. To solve this challenging problem, this paper is dedicated to exploring the potential of solely using labeled clean-weather LiDAR data for training, with a focus on enhancing the Domain Generalization (DG) capabilities of multi-agent perception systems in different unseen adverse weather conditions. We discover and summarize three specific factors for the domain gap between clean to adverse weather conditions of multi-agent perception, as detailed below.
@@ 1 Introduction
  • Perception Range Reduction: Snowflakes, raindrops, or fog particles might temporarily block the LiDAR’s line of sight, causing loss or erroneous data in the point cloud. This dynamic occlusion reduces the vehicle’s ability to perceive its surroundings accurately in the real world.
@@
  • Point Cloud Degradation: Severe weather conditions, such as fog and rain, might significantly change the characteristics compared to clean weather, such as optical attenuation in the atmosphere, surface reflectivity changes, dynamic occlusions. These factors reduce the amount of effective signal returned to the LiDAR, affecting the integrity and accuracy of the point cloud.
@@
  • Accumulated Damages on Perception: The adverse weather conditions impair each agent’s perception, so the multi-agent perception system might have accumulated damages due to the V2X information sharing.
@@
Figure 1: Vehicle-to-Vehicle (V2V) and Vehicle-to-Everything (V2X) examples for multi-agent perception system under adverse weather. Reduced perception range can be found and the point cloud degradation is enlarged. Different-color agent in the left snapshot has the corresponding color in the right point clouds. CAV: Connected and Automated Vehicle, Infra: Infrastructure. Best view in color.
Based on the previous related works [40, 10, 14, 9], there are reduced perception range and point cloud degradation in fog, rain, and snow senarios, which might result in potential accumulated damages to the whole perception system. The domain gaps by the above three factors could result in significant performance decline in multi-agent perception under adverse weather conditions, which have never been well studied before. Due to the difficulties of collecting multi-agent data with communication and considering the adverse weather in the real world, to facilitate research in this area, we leverage the OPV2V [48] and V2XSet [47] datasets of multi-agent perception, by simulating the effects of three common adverse weather conditions, i.e., Fog, Rain, Snow to generate two new adverse weather datasets, namely OPV2V-w and V2XSet-w, to study this scientific problem as shown in Fig. 1.
This paper proposes a novel domain generalization approach called V2X-DGW for LiDAR-based 3D object detection of V2X multi-agent perception systems under adverse weather conditions. This approach is designed to improve performance in unseen adverse weather conditions, training exclusively on the clean-weather source data. First of all, we design a new Adaptive Weather Augmentation (AWA) to mimic the LiDAR point clouds in adverse weather conditions with reduced perception range and diverse degradation. We then propose two alignment techniques for the generalizable representation learning, namely Trust-region Weather-invariant Alignment (TWA) and Agent-aware Contrastive Alignment (ACA). Specifically, TWA is proposed for domain generalization in adverse weather conditions and ACA is used as a contrastive learning based regularization to shrink the accumulated damages on perception due to adverse weather. Finally, we conducted extensive experiments on our two new simulated adverse weather datasets, namely simulated OPV2V-w and V2XSet-w from [48, 47], to justify the effectiveness of our proposed method. Our contributions are summarized as follows.
  • To the best of our knowledge, we propose the first research of domain generalization for LiDAR-based V2X multi-agent perception systems under adverse weather conditions.
  • To assess the impact of adverse weather on multi-agent perception systems, we simulate the effect of three common types of weather conditions: Fog, Rain, and Snow, on point cloud from the clean weather, utilizing the existing OPV2V and V2XSet datasets to create two novel adverse weather benchmark datasets, namely OPV2V-w and V2XSet-w.
@@ 2 Related Work
Multi-agent Perception. Multi-agent perception systems transcend the constraints of single-vehicle systems by utilizing information via collaboration modules, enhancing effectiveness and results of multi-agent perception tasks [43, 48, 39, 47]. Generally, three methods are mainly used for combining observations from multiple vehicles: merging raw data [4, 35], integrating features during processing [22, 3, 39, 48, 20], and output fusion [50, 1]. Attfuse [48] utilizes self-attention models to enhance the sharing features for V2V perception system. SCOPE [49] presents a learning-based framework that targets multi-agent challenges, prioritizing the temporal context of the ego agent. CoBEVT [44] and V2X-ViT [47] introduces a unified Vision Transformer architecture to integrate features from multi-agent perception systems. While these methods demonstrate remarkable performance in multi-agent perception, they have primarily been evaluated under clear weather conditions, neglecting the impact of adverse weather. This paper aims to address these domain gaps by considering the effect of adverse weather conditions on multi-agent perception system.
Deployment of Multi-agent System. While multi-agent perception systems present numerous benefits, establishing a practical real-world multi-agent perception system encounters several obstacles. These challenges include susceptibility to adversarial attacks [20, 37], lossy communication [23, 51], localization inaccuracies [47, 29], communication delay [47], data discrepancy [22, 21], and model heterogeneity [42]. To prevent the erosion of collaborative advantages posed by these challenges, some strategies have been presented to bolster system robustness. S2R-ViT [22] is introduced to address the deployment gap and feature gap between simulation and reality in multi-agent cooperative perception tasks. FDA [21] is proposed to enable cross-domain learning in multi-agent perception systems to reduce data silos caused by distribution gap in multi-agent perception. LCRN [23] mitigates the lossy communication challenge by tackling packet loss during communications. HEAL [28] is introduced as the first extensible heterogeneous collaborative perception framework, establishing a unified feature space and aligning new agents to it. However, adverse weather conditions such as rain, snow, and fog pose significant challenges for single-agent systems relying on camera or LiDAR sensory data [33, 12, 24, 10, 14, 9]. To mitigate this issue, various physics-based simulation methods [14, 9] have been introduced to generate point clouds across diverse weather scenarios. Nevertheless, the impact of adverse weather conditions on multi-agent perception systems has not been previously investigated. Thus, our aim is to enhance the domain generalization capability of multi-agent systems under adverse weather conditions.
@@
Domain Generalization.  Domain generalization (DG) is aim at achieving the excellent performance in unseen domains by exclusively learning from source domains. Current DG methods have been extensively explored object detection [38, 17], semantic segmentation [36, 25, 15, 41] and reinforcement learning [11, 32] tasks in 2D and 3D computer vision tasks. As for strategies of DG employing multiple sources, which encompasses techniques such as the the disentanglement of domain-specific and domain-invariant features [27, 52, 2], alignment of source domain distributions [5, 31, 19], and domain augmentation [55, 54]. For example, INB-Net [34] merges Instance Normalization (IN) with batch normalization to learn appearance invariant feature while maintaining content information. MLDG [18] is introduced as a meta-learning approach that synthesizes virtual testing domains during training to improve generalization to novel domains. These approaches have been instrumental in navigating the complexities associated with learning from diverse source domains. However, the extension of DG to the 3D point cloud domain, particularly concerning LiDAR point clouds, remains relatively untapped because of distinct type of gap stemming from 3D point clouds compared with 2D images. Some recent studies [53, 41, 15] have analyzed the sparsity of point clouds in addressing DG issues. In this paper, we focus on investigating the impact of LiDAR-based object detection on multi-agent perception systems in adverse weather conditions. To reach this goal, we propose a DG approach tailored for multi-agent perception systems operating under adverse weather conditions.
3 Methodology
3.1 Adverse-Weather Benchmark for Multi-agent Perception
Existing datasets on multi-agent perception are all collected under the assumption of clean weather. In this paper, we employ state-of-the-art physics-based simulation methods on the existing multi-agent perception datasets (OPV2V [48] and V2Xset [47]) to create two new benchmarks: OPV2V-w and V2Xset-w, to assess the impact of harsh weather on cooperative perception.
Physics-based Simulation.
We consider three different harsh weather conditions: Fog, Rain, Snow, as they are more common in the real world. Physics-based point cloud simulation methods have shown its success to approximate real weather conditions. Here, we adopt three state-of-the-art physics-based point cloud simulation methods [10, 14, 9] for simulating Fog [^1] https://github.com/MartinHahner/LiDAR_fog_sim, Rain [^2] https://github.com/velatkilic/LISA and Snow [^3] https://github.com/SysCV/LiDAR_snow_sim conditions respectively with the publicized source code. These simulation methods are based on the physical reflection and geometrical optical model in LiDAR sensors. The most recent study [6] have corroborated the consistency of perception performance in synthetic weather data by the physics-based simulation with that observed in real-world data under adverse weather scenarios.
OPV2V-w and V2XSet-w. OPV2V [48] and V2XSet [47] are two large scale open benchmark datasets for LiDAR based multi-agent perception. Since both of them are collected from the same game engine software Carla [7], leveraging these datasets to focus on the impact of weather variations later. Here we deploy physics-based simulation methods on the testing sets of OPV2V and V2XSet, then obtain OPV2V-w and V2XSet-w to evaluate domain generalization under adverse weather conditions.
3.2 Overview of Architecture
The proposed framework of V2X-DGW is depicted in Fig. 3. Given a spatial graph of CAVs within the communication range in clean weather source data, the point clouds of ego and CAVs are denoted $\mathbf{P}^{s}_{ego}\in\mathbb{R}^{4\times m}$ and $\mathbf{P}^{s}_{cav}\in\mathbb{R}^{4\times m}$, respectively. We firstly deploy Adaptive Weather Augmentation $\mathbf{A}_{w}$ on the clean weather data and obtain the augmented data $\mathbf{P}^{a}_{ego}\in\mathbb{R}^{4\times m}$ and $\mathbf{P}^{a}_{cav}\in\mathbb{R}^{4\times m}$. Then all of them are projected into 2D pseudo-image shape $\mathbf{I}\in\mathbb{R}^{C\times H\times W}$ by leveraging the PointPillar [16]’s Pillar Feature Net (PFN), defined as $\mathbf{M}$. In our framework, each CAV shares the same Encoder for LiDAR feature extraction in all conditions. The ego vehicle acquires the visual features of neighboring CAVs through wireless communication. The intermediate features aggregated from $N$ surrounding CAVs in source and augmented data are denoted as $\mathbf{F}^{s}_{cav}\in\mathbb{R}^{N\times H\times W\times C}$ and $\mathbf{F}^{a}_{cav}\in\mathbb{R}^{N\times H\times W\times C}$, respectively. The ego intermediate features are denoted as $\mathbf{F}^{s}_{ego}\in\mathbb{R}^{1\times H\times W\times C}$ and $\mathbf{F}^{a}_{ego}\in\mathbb{R}^{1\times H\times W\times C}$, respectively. The intermediate features are then fused by the same Feature Fusion Network (FFN). Finally, the fused feature maps are fed into the same prediction header for 3D bounding-box regression and classification. Here we formulate the whole framework $\mathbf{DGW}(\cdot)$ as
$\mathbf{I}^{s}_{cav},\mathbf{I}^{a}_{cav},\mathbf{I}^{s}_{ego},\mathbf{I}^{a}_{ego}=\mathbf{M}(\mathbf{P}^{s}_{cav},\mathbf{A}_{w}(\mathbf{P}^{s}_{cav}),\mathbf{P}^{s}_{ego},\mathbf{A}_{w}(\mathbf{P}^{s}_{ego})),$ (1)
$\mathbf{F}^{s}_{cav},\mathbf{F}^{a}_{cav},\mathbf{F}^{s}_{ego},\mathbf{F}^{a}_{ego}=\mathbf{E}(\mathbf{TWA}(\mathbf{I}^{s}_{cav},\mathbf{I}^{a}_{cav}),\mathbf{TWA}(\mathbf{I}^{s}_{ego},\mathbf{I}^{a}_{ego})),$ (2)
$\hat{\mathbf{F}}^{s},\hat{\mathbf{F}}^{a}=\text{FFN}(\mathbf{ACA_{a}}(\mathbf{F}^{s}_{cav},\mathbf{F}^{a}_{cav},\mathbf{F}^{s}_{ego},\mathbf{F}^{a}_{ego})),$ (3)
$\mathbf{DGW}(\mathbf{P}^{s}_{cav},\mathbf{P}^{s}_{ego})=\mathbf{H}(\mathbf{ACA_{g}}(\mathbf{TWA}(\hat{\mathbf{F}}^{s},\hat{\mathbf{F}}^{a}),*)),$ (4)
where $\mathbf{E}(\cdot)$ is the Encoder for LiDAR feature extraction, $\text{FFN}(\cdot)$ is the Feature Fusion Network responsible for fusing the features of CAVs and the ego vehicle, $\hat{\mathbf{F}}$ is the fused feature extracted from $\text{FFN}(\cdot)$ and $\mathbf{H}$ is the prediction header for 3D object detection, $\mathbf{TWA}(\cdot)$ is the proposed Trust-region Weather-invariant Alignment (TWA), $\mathbf{ACA_{a}}(\cdot)$ is the proposed Agent-aware Contrastive Alignment (ACA) in the agent-level, $\mathbf{ACA_{g}}(\cdot)$ is the proposed Agent-aware Contrastive Alignment (ACA) in the group-level, and *** indicates the fused features of other samples in the same batch during training.
@@ 3.2 Overview of Architecture
3.3 Adaptive Weather Augmentation
Based on our above discovery, the adverse weather conditions have 1) perception range reduction and 2) point cloud degradation in the LiDAR based point cloud data. Therefore, we design Adaptive Weather Augmentation (AWA) from two perspectives.

### v5

@@ Abstract
Current LiDAR-based Vehicle-to-Everything (V2X) multi-agent perception systems have shown the significant success on 3D object detection. While these models perform well in the trained clean weather, they struggle in unseen adverse weather conditions with the domain gap. In this paper, we propose a Domain Generalization based approach, named V2X-DGW, for LiDAR-based 3D object detection on multi-agent perception system under adverse weather conditions. Our research aims to not only maintain favorable multi-agent performance in the clean weather but also promote the performance in the unseen adverse weather conditions by learning only on the clean weather data. To realize the Domain Generalization, we first introduce the Adaptive Weather Augmentation (AWA) to mimic the unseen adverse weather conditions, and then propose two alignments for generalizable representation learning: Trust-region Weather-invariant Alignment (TWA) and Agent-aware Contrastive Alignment (ACA). To evaluate this research, we add Fog, Rain, Snow conditions on two publicized multi-agent datasets based on physics-based models, resulting in two new datasets: OPV2V-w and V2XSet-w. Extensive experiments demonstrate that our V2X-DGW achieved significant improvements in the unseen adverse weathers.
I Introduction
The advent of multi-agent perception systems has marked a significant leap forward in surmounting the inherent limitations of single-agent perception from the challenges of perceiving range and occlusion [1, 2]. This paper uses the Connected and Automated Vehicles (CAV) based cooperative perception in autonomous driving as an example of the multi-agent perception systems. Leveraging the Vehicle-to-Everything (V2X) communication for the exchange of LiDAR based point cloud data, these multi-agent systems have demonstrated a substantial enhancement in 3D object detection over the single-agent perception systems [3, 4, 5, 6]. However, the LiDAR-based V2X multi-agent perception systems under the adverse weather conditions (Fog, Rain, Snow) have not been studied and evaluated before.
So far as we know, all the current LiDAR-based V2X multi-agent perception models [2, 7, 5] are trained on the clean-weather data. In our experiments on the publicized V2X cooperative perception dataset OPV2V [4], if we directly deploy these models (trained on clean weather) to the unseen adverse weather conditions, their performance will drop significantly, for example, AP@IoU=0.5 from 88.10 in clean weather to 67.55 in fog weather. We discover and summarize three specific challenges for the domain gap between clean and adverse weather conditions of multi-agent perception, as detailed below.
@@ I Introduction
  • Perception Range Reduction: As shown in Fig. 1, the LiDAR’s perception range is reduced in the adverse weather. The fog particles, raindrops, snowflakes occasionally block the LiDAR’s line of sight causing lossy data in the point cloud, so this dynamic occlusion reduces the vehicle’s ability to perceive its surroundings accurately.
@@
  • Point Cloud Degradation: As highlighted by the blue arrows in Fig. 1, the LiDAR’s point cloud quality is degraded. Severe weather conditions might change the characteristics compared to clean weather, such as optical attenuation in the atmosphere, surface reflectivity changes, dynamic occlusions. These factors reduce the amount of effective signal returned to the LiDAR, affecting the integrity and accuracy of the point cloud.
@@
  • Accumulated Damages on Perception: Intuitively, the adverse weather conditions impair each agent’s perception, so the multi-agent perception system might have accumulated damages due to the agent-to-agent information sharing.
@@
Figure 1: Multi-agent perception system under adverse weather. Three CAVs’ LiDAR point clouds are highlighted in Black, Orange and Green colors. We can discover the reduced perception range (shrinked wave circles) and degradation (highlighted by the blue arrows) for the point clouds under adverse weather.
To solve the above three challenges, this paper proposes a novel Domain Generalization (DG) approach called V2X-DGW for LiDAR-based 3D object detection of V2X multi-agent perception under adverse weather conditions. Our approach is designed to improve performance in unseen adverse weather conditions, training exclusively on the clean-weather source data. First of all, we design a new Adaptive Weather Augmentation (AWA) to mimic the LiDAR point clouds in adverse weather conditions with reduced perception range and diverse degradation. We then propose two alignment techniques for the generalizable representation learning, namely Trust-region Weather-invariant Alignment (TWA) and Agent-aware Contrastive Alignment (ACA). Specifically, TWA is proposed for domain generalization in the unseen adverse weather conditions and ACA is used as a contrastive learning based regularization to shrink the accumulated damages on perception due to adverse weather.
Furthermore, due to the difficulties of collecting multi-agent data with communication under adverse weathers, there are currently no publicized benchmark datasets to investigate this research problem. To facilitate this research, we leverage the publicized OPV2V [4] and V2XSet [8] datasets of V2X based multi-agent perception, by simulating three common adverse weather conditions (Fog, Rain, Snow) by the physics-based point cloud models [9, 10, 11] to generate two new adverse weather V2X datasets, namely OPV2V-w and V2XSet-w. Finally, we conducted extensive experiments on our simulated OPV2V-w and V2XSet-w datasets, to justify the effectiveness of our proposed method. Our contributions are summarized as follows.
  • To the best of our knowledge, we propose the first research of Domain Generalization for LiDAR-based V2X multi-agent perception systems under unseen adverse weather conditions.
  • To assess the impact of adverse weather on multi-agent perception systems, we simulate the effect of three common types of weather conditions(Fog, Rain, Snow), on point cloud from the clean weather, utilizing the existing OPV2V and V2XSet datasets to create two new adverse weather benchmark datasets, namely OPV2V-w and V2XSet-w.
@@ II Related Work
Multi-agent Perception. Multi-agent perception systems transcend the constraints of single-vehicle systems by utilizing information via collaboration modules, enhancing effectiveness and results of multi-agent perception tasks [12, 4, 3, 8]. Generally, three methods are mainly used for combining observations from multiple vehicles: merging raw data [13], integrating features during processing [5, 1, 3, 4, 6], and output fusion [14, 15]. Attfuse [4] utilizes self-attention models to enhance the sharing features for V2V perception system. SCOPE [16] presents a learning-based framework that targets multi-agent challenges, prioritizing the temporal context of the ego agent. V2X-ViT [8] introduces a unified Vision Transformer architecture to integrate features from multi-agent perception systems. While these methods demonstrate remarkable performance in multi-agent perception, they have primarily been evaluated under clear weather conditions, neglecting the impact of adverse weather.
@@
Domain Generalization.  Domain generalization (DG) aims at achieving the excellent performance in unseen domains by exclusively learning from source domains. Current DG methods have been extensively explored on object detection [17, 18], semantic segmentation [19, 20, 21, 22] and reinforcement learning [23, 24] tasks in 2D and 3D computer vision tasks. As for strategies of DG employing multiple sources, which encompasses techniques such as the the disentanglement of domain-specific and domain-invariant features [25, 26, 27], alignment of source domain distributions [28, 29, 30], and domain augmentation [31]. These approaches have been instrumental in navigating the complexities associated with learning from diverse source domains. Some recent studies [32, 22, 21] have analyzed the sparsity of point clouds in addressing DG issues. However, the extension of DG to the LiDAR-based 3D object detection on multi-agent perception systems in adverse weather conditions remains relatively untapped. This paper aims to design a new DG method to promote the LiDAR-based 3D object detection on multi-agent perception systems in adverse weather conditions.
III Methodology
@@
III-A Adverse-Weather Benchmark for Multi-agent Perception
Existing multi-agent perception datasets are all collected under the assumption of clean weather. We employ state-of-the-art physics-based simulation methods on the existing multi-agent perception datasets (OPV2V [4] and V2Xset [8]) to create two new benchmarks: OPV2V-w and V2Xset-w, to assess the impact of harsh weather.
Physics-based Simulation. We consider three different harsh weather conditions: Fog, Rain, Snow, as they are common in the real world. Physics-based point cloud simulation methods have shown its success to approximate real weather conditions. We adopt three state-of-the-art physics-based point cloud simulation methods [9, 10, 11] for simulating Fog [^1] https://github.com/MartinHahner/LiDAR_fog_sim, Rain [^2] https://github.com/velatkilic/LISA and Snow [^3] https://github.com/SysCV/LiDAR_snow_sim conditions respectively with the publicized source code. These simulation methods are based on the physical reflection and geometrical optical model in LiDAR sensors. The most recent study [34] have corroborated the consistency of perception performance in synthetic weather data by the physics-based simulation with that observed in real-world data under adverse weather scenarios.
OPV2V-w and V2XSet-w. OPV2V [4] and V2XSet [8] are two large scale datasets for LiDAR based multi-agent perception. Since both of them are collected from the same game engine software Carla [35], leveraging these datasets to focus on the impact of weather variations later. We deploy physics-based simulation methods on the testing sets of OPV2V and V2XSet, then obtain OPV2V-w and V2XSet-w to evaluate domain generalization under adverse weather.
III-B Overview of Architecture
((a))
Figure 3: Illustration of Adaptive Weather Augmentation.
((a))
Our V2X-DGW is depicted in Fig. 2. Given a spatial graph of CAVs within the communication range in clean weather source data, the point clouds of ego and CAVs are denoted $\mathbf{P}^{s}_{ego}\in\mathbb{R}^{4\times m}$ and $\mathbf{P}^{s}_{cav}\in\mathbb{R}^{4\times m}$, respectively. We firstly deploy Adaptive Weather Augmentation $\mathbf{A}_{w}$ on the clean weather data and obtain the augmented data $\mathbf{P}^{a}_{ego}\in\mathbb{R}^{4\times m}$ and $\mathbf{P}^{a}_{cav}\in\mathbb{R}^{4\times m}$. Then all of them are projected into 2D pseudo-image shape $\mathbf{I}\in\mathbb{R}^{C\times H\times W}$ by leveraging the PointPillar [33]’s Pillar Feature Net (PFN), defined as $\mathbf{M}$. So, $\mathbf{I}^{s}_{cav},\mathbf{I}^{a}_{cav},\mathbf{I}^{s}_{ego},\mathbf{I}^{a}_{ego}=\mathbf{M}(\mathbf{P}^{s}_{cav},\mathbf{A}_{w}(\mathbf{P}^{s}_{cav}),\mathbf{P}^{s}_{ego},\mathbf{A}_{w}(\mathbf{P}^{s}_{ego})).$ In our framework, each CAV shares the same Encoder for LiDAR feature extraction in all conditions. The ego vehicle acquires the visual features of neighboring CAVs through wireless communication. The intermediate features aggregated from $N$ surrounding CAVs in source and augmented data are denoted as $\mathbf{F}^{s}_{cav}\in\mathbb{R}^{N\times H\times W\times C}$ and $\mathbf{F}^{a}_{cav}\in\mathbb{R}^{N\times H\times W\times C}$, respectively. The ego intermediate features are denoted as $\mathbf{F}^{s}_{ego}\in\mathbb{R}^{1\times H\times W\times C}$ and $\mathbf{F}^{a}_{ego}\in\mathbb{R}^{1\times H\times W\times C}$, respectively. The intermediate features are then fused by the same Feature Fusion Network (FFN). Finally, the fused feature maps are fed into the same prediction header for 3D bounding-box regression and classification. Here we formulate the whole framework $\mathbf{DGW}(\cdot)$ as
$\mathbf{F}^{s}_{cav},\mathbf{F}^{a}_{cav},\mathbf{F}^{s}_{ego},\mathbf{F}^{a}_{ego}=\mathbf{E}(\mathbf{T}_{w}(\mathbf{I}^{s}_{cav},\mathbf{I}^{a}_{cav}),\mathbf{T}_{w}(\mathbf{I}^{s}_{ego},\mathbf{I}^{a}_{ego})),$ (1)
$\hat{\mathbf{F}}^{s},\hat{\mathbf{F}}^{a}=\text{FFN}(\mathbf{C}_{a}(\mathbf{F}^{s}_{cav},\mathbf{F}^{a}_{cav},\mathbf{F}^{s}_{ego},\mathbf{F}^{a}_{ego})),$ (2)
$\mathbf{DGW}(\mathbf{P}^{s}_{cav},\mathbf{P}^{s}_{ego})=\mathbf{H}(\mathbf{C}_{g}(\mathbf{T}_{w}(\hat{\mathbf{F}}^{s},\hat{\mathbf{F}}^{a}),*)),$ (3)
where $\mathbf{E}(\cdot)$ is the Encoder for LiDAR feature extraction, $\text{FFN}(\cdot)$ is the Feature Fusion Network responsible for fusing the features of CAVs and the ego vehicle, $\hat{\mathbf{F}}$ is the fused feature extracted from $\text{FFN}(\cdot)$ and $\mathbf{H}$ is the prediction header for 3D object detection, $\mathbf{T}_{w}(\cdot)$ is the proposed Trust-region Weather-invariant Alignment (TWA), $\mathbf{C}_{a}(\cdot)$ is the proposed Agent-aware Contrastive Alignment (ACA) in the agent-level, $\mathbf{C}_{g}(\cdot)$ is the proposed Agent-aware Contrastive Alignment (ACA) in the group-level, and * indicates the fused features of other samples in the same batch.
III-C Adaptive Weather Augmentation
Based on our above discovery, the adverse weather conditions have 1) perception range reduction and 2) point cloud degradation in the LiDAR based point cloud data. We design Adaptive Weather Augmentation (AWA) from two views.